An express train and a local train both leave Gray’s Lake at 12:00 noon and head for Chicago 60 miles away. The express travels twice as fast as the local and arrives 2 hours ahead of it. Find the speed of each train.
Answer by checkley77(12844)

D=RT
60=2R(T-2)
60=R(T+2)
60=2RT-4R
60=RT+2R MULTIPLY THIS EQUATION BY -2 & ADD
-120=-2RT-4R
-60=2RT-4R
------------------
-60=-8R
-8R=-60
R=-60/-8
R=7.5 MPH FOR THE LOCAL TRAIN.
2*7.5=15 MPH FOR THE EXPRESS TRAIN.
60=15(T-2)
60=15T-30
15T=90
T=90/15
T=6 HOURS FOR THE EXPRESS
60=7.5(6+2)
60=7.5*8
60=60